Question 3 (15 scores) Groucho, Harpo, and Chico go into partnership on January 1, 2015. Groucho contributes Harpo $70,000, and Chico $40,000 to a business called Marx Brothers Partnership a monthly basis, each partner is allocated income and is allowed to receive cash from the business in proportion to the capital they prov ided. Assume that Groucho receives $2,700 cash per month. Required: Part Prepare the journal entry for the initial investment. a. Part b. Det c. Prepare the journal entry that would be made in one month for the monthly distribution. ermine the monthly distribution amounts for each of the three partners Part d. Prepare the journal entry for the allocation of an annual net income of $84,000. For purposes of this journal entry, assume Sales Revenue totaled $116,000 and that all expenses, totaling $32,000, were recorded in a single account called Operating Expenses. Part e. Prepare the journal entry to close the Drawings accounts at the end of the year. Part f. Prepare a Statement of Partners Equity (assume no additional investments made)
Answer #1

Part a.

In the books of Marx Brothers Partnership :

Date Account Titles Debit Credit
$ $
January 1, 2015 Cash 200,000
Groucho, Capital 90,000
Harpo, Capital 70,000
Chicco, Capital 40,000

Part b:

Capital contribution ratio :

Groucho : 90,000 / 200,000 = 0.45

Harpo : 70,000 / 200,000 = 0.35

Chicco : 40,000 / 200,000 = 0.20

If Groucho receives $ 2,700 per month, total cash distribution = $ 2,700 / 0.45 = $ 6,000.

Harpo will receive $ 6,000 x 0.35 = $ 2,100.

Chicco will receive $ 6,000 x 0.20 = $ 1,200

Part c :

Date Account Titles Debit Credit
$ $
January 31, 2015 Groucho, Drawings 2,700
Harpo, Drawings 2,100
Chico, Drawings 1,200
Cash 6,000

Part d :

Date Account Titles Debit Credit
$ $
December 31, 2015 Sales Revenue 116,000
Groucho, Capital ( $ 84,000 x 0.45) 37,800
Harpo, Capital ( $ 84,000 x 0.35) 29,400
Chicco, Capital ( $ 84,000 x 20) 16,800
Operating Expenses 32,000

Part e.

Date Account Titles Debit Credit
$ $
December 31, 2015 Groucho, Capital 32,400
Groucho, Drawings 32,400
December 31, 2015 Harpo, Capital 25,200
Harpo, Drawings 25,200
December 31, 2015 Chicco, Capital 14,400
Chicco, Drawings 14,400

Part f:

Marx Brothers' Partnership
Statement of Partners' Equity
For the year ended December 31, 2015
Groucho Harpo Chico
Capital contribution , January 1, 2015 $ 90,000 $ 70,000 $ 40,000
Add: Share of Net Income 37,800 29,400 16,800
Total 127,800 99,400 56,800
Less: Drawings (32,400) (25,200) (14,400)
Partners' Equity, December 31, 2015 $ 95,400 $ 74,200 $ 42,400